# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v1

@@ Abstract
Despite considerable efforts on making them robust, real-world ML-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable. The canonical approach in robustness evaluation calls for adaptive attacks, that is with complete knowledge of the defense and tailored to bypass it. In this study, we introduce a more expansive notion of being adaptive and show how attacks but also defenses can benefit by it and by learning from each other through interaction.
@@
We validate our observations through a wide theoretical and empirical investigation to confirm that AI-enabled adversaries pose a considerable threat to black-box ML-based systems, rekindling the proverbial arms race where defenses have to be AI-enabled too. Succinctly, we address the challenges posed by adaptive adversaries and develop adaptive defenses, thereby laying out effective strategies in ensuring the robustness of ML-based systems deployed in the real-world.
@@ I Introduction
In this study we address a crucial gap, as in adversarial machine learning (AML) evaluating the robustness of defenses against oblivious, non-adaptive, and thus suboptimal attackers is inherently problematic [50, 19]. We expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adaptively control their parameters and evasive actions together in response to how the model under attack and its defenses respond [5]. We demonstrate theoretically and empirically how self-adaptive adversaries can modify their own policies through reinforcement learning (RL) to become both optimal and evade active detection. Notably, we find that this can be performed in a gradient-based manner even in fully black-box contexts [III.1], a capability that properly reflects the level of adversarial threat and does not overestimate the empirical robustness; attackers might compute gradients after all.
@@
This capability, however, enables active defenses in turn, as through rigorous threat modeling and by simulating self-adaptive attackers their full potential is uncovered and effective counter-policies can be learned. To express the call for adaptive evaluations in AML differently, a defense can only be as good as its adversary. This mutual interdependence generates the necessity for both attacks and defenses being self-adaptive, as well as the competitive, zero-sum game they jointly form. In summary, our research engages from two related perspectives: a) countering decision-based attacks against ML models, and b) the self-adaptation of such attacks, resulting in the following contributions:
@@
  • We demonstrate that active defenses against decision-based attacks are a necessary complement to model hardening but not sufficient. Active defenses are inevitably bypassed by self-adaptive attackers; and in turn this necessitates self-adaptive defenses too.
@@
  • To facilitate reasoning on self-adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG). We demonstrate how attacks can optimize their policy and evade active detection at the same time; as a counter we propose a novel active defense and employ RL agents to adapt and optimize both.
@@
  • In an extensive empirical evaluation on image classification and across various adversarial settings, we validate our theoretical analysis and show that self-adaptation through RL outperforms baseline attacks, evasion techniques, and model hardening defenses like adversarial training.
In the domain of black-box, decision-based AML, our work indicates that robust evaluations cannot consider attacks or defenses in isolation, and that each should have the capability to modify their operation through interaction and in direct response to other agency in the environment.
II Related Work & Problem Statement
@@
In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are regarded as a highly realistic and pervasive threat in AI-based cybersecurity environments. Despite the lack of the closed-form expression of the model under attack, given enough queries the effectiveness of such black-box attacks can match and even surpass that of white-box techniques like C&W [14].
@@ II-A Attacks & Mitigations
The partial exception to this rule is adversarial training [35]. Given dataset $D={(x_{i},y_{i})}^{n}_{i=1}$ with classes $C$ where $x_{i}\in\mathbb{R}^{d}$ is a clean example and $y_{i}\in{1,...,C}$ is the associated label, the objective of adversarial training is to solve the following min-max optimization problem:
@@ II-B Stateful Defenses
All decision-based attacks share properties that can be useful in devising defenses against them, beyond adversarial training. Such a property is their inherent sequentiality: as they follow a query generation policy towards the optimal adversarial example, the generated samples are correlated. This might not be the case for the queries themselves however, as the adversary might adaptively employ transformations the model is invariant to, like the query blinding strategy in [16]. This work is also the first to employ a stateful defense against query-based attacks. Another one is PRADA [30], a defense built against model extraction but which also works against model evasion. The efficacy of these approaches however rests on the assumption that queries can be consistently linked to uniquely identifiable actors – that also show limited to no collaboration – so that a buffer of queries can be built for each one of them.
This limitation was recently addressed, together with the ensuing scalability issues in the Blacklight defense, by resourcefully employing hashing and quantization [32]. It remains a similarity/fingerprinting defense however, thus vulnerable to circumvention if an adversary can find a query generation policy to preserve the attack functionality while evading detection; this was recently achieved by adapting existing attacks through the rejection signal that Blacklight returns [21]. Ultimately, any stateful defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be representative only if the adversary has exhausted their adaptive offensive capabilities.
From the mere interaction it is enabled from, stateful detection already points towards adversarial examples being something broader that includes the process of generating them, otherwise defined as adversarial behavior. Furthermore, as AML encompasses domains where visual similarity is of less significance [45], the ensuing question is where does the boundary lie exactly between benign and adversarial examples? In AI-based environments where adversarial activity is also present, we posit that adversarial behavior is a more general formulation, appropriate for describing and countering threats at the interface of interaction. It is also more intuitive in domains where behavior – as information unfolding over time – is the canonical description, e.g in malware or bot detection, where AI can imitate benign programs or human behavior respectively.
@@ II-C On Being Adaptive
The correct way to evaluate any proposed defense is against adaptive attacks: that is, attacks with explicit knowledge of the inner mechanisms of a defense [50]. This has been established in computer security as the stipulation that security through obscurity does not work: every defense has a potential workaround. If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, an active defense like stateful detection is the counterpart to decision-based attacks, and as we will demonstrate further in the paper, also the necessary complement to model hardening.
At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated. As demonstrated in the work of Croce et al. [19], it is very common that the parameters of an attack are suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness. This underestimation can be further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have. The practical effectiveness of attacks rests on the ability to adapt the policies that govern both their operation and their evasive capabilities in tandem.
In AML, “adaptive” refers conventionally to attacks that have full knowledge of the inner workings of a defense; a term we will henceforth refer to as adapted. In our work, we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [5].
Typically, what is to be controlled is known in advance and well-defined. The moment however we consider adaptive evaluations, new controls are immediately implied: in a newly proposed defense, for instance rejection-based, these are controls that can bypass it, like rejection sampling. To flesh out the twofold meaning of adaptive, one has to both imagine new knobs [29], and discover their correct configuration that would lead to the optimal result. We conceptualize this expanded definition of adaptive, essential for having accurate evaluations in AML research, in Figure 1.
Figure 1: In AML adaptive means to invent new knobs that can bypass a defense (lower open set); in control theory it means the precise tuning of all known knobs. In this work, we reformulate adaptive to signify both. In HSJA for example, radius, steps, and jumps are parameters of the attack, while rotate and translate are evasive transformations.
@@ II-D Research Gap
Prior work has focused on adapted attacks by incorporating general knowledge of any defenses and then empirically configured to evade it [14, 15, 9]. Defenses also follow the same adapted paradigm, with parameters empirically defined and then fixed for deployment [16, 32]. Our observation here is that neither attacks nor defenses are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and against other adaptive agents in it, something that comes with considerable limitations in security applications where the latter is a given. To bridge this gap, we provide theoretical treatment and empirical study of existing and novel methodologies adapting through direct interaction with their environment, and denote them as self-adaptive.
Our work builds on a long line of prior research that focuses on different aspects of the competition between adversaries and defenses. Carlini and Wagner [14] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation. Bose et al[8] propose Adversarial Examples Games (AEG), a zero-sum game between a white-box attacker and a local surrogate of the target model family. At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples. Pal et al[37] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium. Ren et al[42] explain adversarial robustness through low and high-order interactions between features. Yasodharan et al[56] study adversarial classification by formulating it as non-zero-sum hypothesis testing game.
As the most relevant and representative setting for threats against deployed AI-based systems, we scope on decision-based, interactive attacks and defenses; we contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive (and thus also the defensive) potential is properly assessed. While “self-adaptive” represents the synthesis of adapted and adaptive control, for the remainder of the paper we employ the term “adaptive” interchangeably, and we use “adapted” to indicate what is conventionally known as adaptive evaluations in AML.
@@ III Theoretical Framework
In this work we engage from two perspectives that are interrelated: a) thwarting decision-based attacks, and b) adapting such attacks together with any evasive actions, as these are in a natural trade-off. Evaluating non-adaptive – in the expansive sense we outlined – attacks or defenses renders any results unreliable and incomplete [50]. When offensive or defensive methodologies become adaptive, the environments that they reside in become non-stationary [28], putting further pressure on the IID foundations that ML builds on. We can approach this multi-agent interaction more generally as a competitive and sequential zero-sum game [34, 26, 8]. To understand the implications of attacks and defenses turning adaptive, before the practical implementation we perform a theoretical investigation of their interactions around an ML-based system, and we initiate our analysis from adversaries as the agency generates all the subsequent reasoning. In the following sections, the notation and terms we introduce are highlighted in red.
@@ III-A Attacks
The most compelling threat for deployed ML-based systems are hard-label, decision-based black-box attacks where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard labels. One of the first decision-based attacks was Boundary Attack [9]. A large number of others followed – each inventive in its own way – that manage to improve the overall performance, typically measured as the minimum perturbation achieved for the minimal amount of queries submitted. Prominent examples of such attacks are HSJA [15], Guessing Smart (BAGS) [12], Sign-Opt [17], Policy-driven (PDA) [55], QEBA [33], and SurFree [36].
White-box methodologies for constructing adversarial examples like C&W [14] cannot function in black-box environments where there is no closed-form description of the inference pipeline. To facilitate optimization, decision-based attacks commonly reverse the adversarial task by initiating from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation. The task then regains the ability to be solved continuously as it allows for the minimization of the perturbation while staying on the adversarial side of the boundary. Decision-based attacks exhibit further commonalities in their functioning so that one can abstract through their particularities: given starting and original samples $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{g}$ and $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{c}$ respectively, the goal is to iteratively find adversarial candidates $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{t}$, until the distance $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\delta=d(x_{t},x_{c})$ is minimized. This process can follow different algorithmic approaches representing different geometrical intuitions, however to describe them generally we can use a candidate generation policy:
@@
that is conditional on $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack. Such parameters are the source and orthogonal steps for BAGS and the number of queries and radius for the gradient approximation in HSJA. Assuming the model answers always and truthfully, this process can be construed as an Markov Decision Process (MDP) to be solved by a policy that minimizes $d$ in the least amount of queries.
Given a multinomial image classifier $\mathcal{M}$, the model under attack is based on a discriminant function $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes. By definition, black-box environments provide no access to these probabilities; instead one can only observe the result of the classifier $C$ that returns the class with maximum probability:
@@
with $D$ being $\operatorname*{arg\,max}$, the decision function. The goal in targeted attacks is to change the decision $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}c_{g}\in[m]$ for a correctly classified example $x$, to a predefined target class $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}c_{o}\neq c_{g}$. This process can be facilitated through a function $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\psi$ which given a perturbed example $x_{t}$ at step $t$, it returns a binary indicator of success:
@@
As long as the model responds, $\psi$ can always be evaluated, it thus constitutes the essential mechanism upon which decision-based attacks build. The adversarial goal can then be described as the following constrained optimization problem:
@@
where the distance metric $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$. As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this optimization task highlights the imprecise nature of adversarial examples, something that is further aggravated in domains where visual affinity is of little importance. Customarily, successful or unsuccessful adversarial examples are delimited by an threshold $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$. While this constraint might not intuitively translate to non-visual domains, we nonetheless contend that minimal perturbation still remains a central property of adversarial examples.
The most relevant attacks in cybersecurity contexts are black-box and decision-based, and they are becoming increasingly effective. In HSJA for example, its optimization is guaranteed to converge to a stationary point of Eq. (5), and given typical $\epsilon$ values on perturbation imperceptibility this translates to near-perfect attack success rates, even against adversarially trained models. The limitations of adversarial training against decision-based attacks can be attributed to the fundamentally out-of-distribution (OOD) nature of adversarial examples, as that makes the saddle point optimization of Eq. (1) intractable to solve exhaustively. Additionally, it is challenging to incorporate decision-based attacks during stochastic gradient descent: as approaches that navigate the decision boundary, the further the latter is from convergence, the less effective the attack is. Scalability is also an issue as typically in adversarial training a few steps (1-50) of a white-box attack – FGSM or PGD – are required, while decision-based attacks can take orders of magnitude more steps (queries) to produce an adversarial example.
Decision-based attacks typically search for the optimal parameters $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\theta$ of generation policy (2), those that given $x^{i}_{c}$, with $i$ denoting the i-th adversarial episode, minimize Eq. (5) in expectation:

### v2

@@ Abstract
Despite considerable efforts on making them robust, real-world ML-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable. The canonical approach in robustness evaluation calls for adaptive attacks, with complete knowledge of the defense and tailored to bypass it. In this study, we introduce a more expansive notion of being adaptive and show how attacks but also defenses can benefit by it and by learning from each other through interaction.
@@
We validate our observations through a wide theoretical and empirical investigation to confirm that AI-enabled adversaries pose a considerable threat to black-box ML-based systems, rekindling the proverbial arms race where defenses have to be AI-enabled too. Our approach outperforms the state-of-the-art black-box attacks and defenses, while bringing them together to render effective insights over the robustness of real-world deployed ML-based systems.
@@ I Introduction
In this study we address a crucial gap, as in adversarial machine learning (AML) evaluating the robustness of defenses against oblivious, non-adaptive, and therefore suboptimal attackers is inherently problematic [14, 15]. We expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adaptively control their parameters and evasive actions together in response to how the model under attack and its defenses respond [16]. We demonstrate theoretically and empirically how self-adaptive attacks can use reinforcement learning (RL) to modify their policies to become both optimal and evade active detection. Notably, this can be performed in a gradient-based manner even in fully black-box contexts [III.3], and is a capability that properly reflects the level of adversarial threat and in that way does not overestimate the empirical robustness; real attackers might compute gradients after all.
@@
This capability, however, enables active defenses in turn. Through rigorous threat modeling and by simulating self-adaptive attackers, their full potential is uncovered and effective counter-policies can be learned. To express the call for adaptive evaluations in AML differently, a defense is trustworthy only insofar the adversary is optimal. This mutual interdependence creates the imperative for both attacks and defenses to be self-adaptive, thereby establishing the competitive, zero-sum game that their interaction forms. Our research approaches the question of robustness from these two closely related perspectives: first, how to optimize decision-based attacks, and secondly what defensive options exist. In this manner, we explore both offensive and defensive strategies in depth, resulting in the following key contributions:
@@
  • We demonstrate that active defenses against decision-based attacks are a necessary but insufficient complement to model hardening. Active defenses are inevitably bypassed by self-adaptive attackers however, and necessitate self-adaptive defenses too.
@@
  • To facilitate reasoning on adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG). We demonstrate how adversaries can optimize their policy and evade active detection at the same time; as a counter, we propose a novel active defense and employ RL agents to adapt and optimize both. For reproducibility and follow-up work, we open-source our code [^1] https://anonymous.4open.science/r/AMG-AD16.
@@
  • In an extensive empirical evaluation on image classification and across various adversarial settings, we validate our theoretical analysis and show that self-adaptation through RL outperforms vanilla black-box attacks, model hardening defenses like adversarial training, and notably both the state-of-the-art adaptive attacks and stateful defenses.
Our work highlights that in the domain of black-box AML, robust evaluations should go a step further than adapting attacks: both attacks and defenses should have the capability to optimize their operation through interaction and in direct response to other agency in their environment. The remainder of the paper is structured as follows: Section II provides the necessary background on the domain and reviews the related work. Section III introduces and motivates our theoretical analysis of robustness under decision-based attacks. Section IV elucidates the threat model and the concrete design choices. In Section V we elaborate on our experimentation and analyze our results. We conclude with Section VI where we discuss insights, limitations and challenges.
II Preliminaries
@@
In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are regarded as a highly realistic and pervasive threat in AI-based cybersecurity environments. Despite the lack of the closed-form expression of the model under attack, given enough queries the effectiveness of such black-box attacks can match and even surpass that of white-box techniques [17, 15].
@@ II-A Attacks & Mitigations
The partial exception to this rule is adversarial training. Given dataset $D={(x_{i},y_{i})}^{n}_{i=1}$ with classes $C$ where $x_{i}\in\mathbb{R}^{d}$ is a clean example and $y_{i}\in{1,...,C}$ is the associated label, the objective of adversarial training is to solve the following min-max optimization problem:
@@ II-B Stateful Defenses
All decision-based attacks share properties that can be useful in devising defenses against them, on top of adversarial training. Such a property is their inherent sequentiality: by following an attack policy towards the optimal adversarial example, the generated candidates are correlated. This might not be the case for the queries themselves however, as the adversary might employ transformations the model is invariant to, like the query blinding strategy in Chen et al. [13]. This work is also the first to employ a stateful defense against query-based attacks. Another stateful defense is PRADA [22], built against model extraction but which also works against model evasion. The efficacy of these approaches however rests on the assumption that queries can be consistently linked (through metadata like IP or account, cf. LABEL:tab:comparison) to uniquely identifiable actors – that also show limited to no collaboration – so that a buffer of queries can be built for each.
This limitation was recently addressed, together with the ensuing scalability issues in the Blacklight defense, by resourcefully employing hashing and quantization [23]. It remains a similarity-based defense however, thus vulnerable to circumvention if an adversary can find a query generation policy that preserves attack functionality while evading detection; the recent OARS work achieved this by adapting existing attacks through the rejection signal Blacklight returns [24]. Ultimately, any (stateful) defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be representative only if the adversary has exhausted their adaptive offensive capabilities.
Figure 1: In AML, adaptive attacks are those with the capabilities (knobs) to bypass a defense; adaptive control is rather the precise tuning of all the known knobs. Against black-box systems, we can reformulate adaptive so that it signifies both. For instance in HSJA [7], radius, steps, and jump are parameters of the attack, while rotate and translate are transformations that can evade a similarity-based defense.
@@ II-C On Being Adaptive
The correct way to evaluate any proposed defense is against adaptive attacks, that is with explicit knowledge of the inner mechanisms of a defense [14]. In computer security this is known as the stipulation that security through obscurity does not work, as the robustness of defenses should not rely on keeping their way of functioning secret. If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, active defenses like stateful detection are the counterpart to decision-based attacks, and as we will further demonstrate, also the necessary complement to hardening a model against them.
At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated. Previous work has demonstrated that loss functions and parameters of attacks are often suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness [15, 25]. This underestimation is further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have. The practical effectiveness of attacks therefore rests on the ability to adapt the policies that govern their operation and their evasive capabilities in tandem.
In AML, “adaptive” by convention refers to attacks with full knowledge of how a defense works and the tools to bypass it; we denote such attacks as adapted. In our work, we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [16]. One can think of adaptive control in the sense “attack optimization” is used by Pintor et al. [25], but for black-box systems. Typically, what is to be controlled is well-defined and known in advance. The moment however we consider adaptive evaluations, new controls are potentially implied: in a similarity-based defense for instance, such controls would be input transformations the model is invariant to. To flesh out the twofold meaning of adaptive, one has to both invent new knobs [26] (which is the conventional understanding of adaptive, and still very hard to automate), and dynamically control their correct configuration that would lead to the optimal result (self-adaptive). We conceptualize this more general definition of adaptive, essential for having accurate evaluations against decision-based attacks, in Figure 1.
@@ II-D Research Gap
Prior work has focused on adapted attacks, which incorporate general knowledge of any defenses and empirically configured to evade it [17, 7, 6]. Defenses also follow the same adapted paradigm of empirically defined and fixed parameters [13, 23]. Our observation is that neither of them are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and with respect to other adaptive agents in it, with clear limitations when the latter is a given, e.g. in cybersecurity. To bridge this gap, we provide a theoretical treatment and empirical study of existing and novel methodologies adapting through direct interaction with their environment, denoting them as self-adaptive.
Our work builds on a long line of prior research that focuses on both sides of the competition between adversaries and defenses. Carlini and Wagner [17] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation. Bose et al. [27] propose Adversarial Examples Games (AEG), a zero-sum game between a white-box attacker and a local surrogate of the target model family. At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples. Pal et al. [28] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium. Feng et al. [24] introduce OARS: adaptive versions of existing attacks that bypass Blacklight [23], the state-of-the-art stateful defense. To function, OARS presupposes the rejection signal that a defense like Blacklight returns; a strong assumption that as we show in this work does not have to hold for stateful defenses. As we demonstrate in section V and Table III, Blacklight can be bypassed without assuming rejection, while the novel stateful defense we introduce can fully withstand the OARS attack.
As the most relevant and representative threat against real-world AI systems, in this work we scope on decision-based, interactive attacks and defenses. We contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive and thus also the defensive potential is properly assessed. In the remainder of the paper the term “adaptive” subsumes adaptive control, and is used interchangeably with “self-adaptive”. For what is conventionally known as adaptive evaluations in AML, we use the term “adapted”. To facilitate comparison, in LABEL:tab:comparison we highlight the most important aspects of our work as the synthesis of adaptive black-box attacks and defenses in a unified framework, and situate it with respect to other relevant and state-of-the-art works in AML. Note the importance for an attack to function without assuming rejection, and respectively for a defense to function without access to query metadata like IP addresses or accounts.
TABLE I: Prominent decision-based attacks and defenses and their individual aspects. While all works are situated either on the offensive or the defensive side, ours brings these two together.
Work Offensive Defensive
Optimized Evasive Adaptive $\neg$¬Rejection Active Adaptive $\neg$¬Metadata Misdirection
Boundary (2018) [6]
BAGS (2018) [29]
HSJA (2020) [7]
OARS (2023) [24]
Adv. Training (2017) [4]
Stateful (2020) [13]
Blacklight (2022) [23]
AMG – Our work
@@ III Theoretical Framework
To investigate the robustness to evasion of real-world systems, two closely related perspectives are pertinent: a) thwarting decision-based attacks, and b) adapting attacks and evasive capabilities in tandem. For the latter, when attacks (and defenses) are evaluated in a non-adaptive manner, in the expansive sense we outlined in subsection II-C, results are rendered incomplete and untrustworthy [14, 25]. Nevertheless, when offensive or defensive techniques adapt the environments that they inhabit become non-stationary [30], putting further pressure on the IID foundations that ML builds on. To understand the implications of such techniques adapting, we perform a theoretical analysis of the possible interactions around the exposed interface of an ML-based system, interactions that can be more generally considered as sequential zero-sum games [31, 32, 27]. We initiate our analysis from adversaries, as the original factor and cause. In the following sections, introduced terms and notation are highlighted in orange.
@@ III-A Attacks
The most compelling threat that deployed ML-based systems face are decision-based black-box attacks, where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard-label responses. One of the first decision-based attacks was Boundary Attack [6]. A large number of others followed – each inventive in its own way – that manage to improve the overall performance, typically measured as the lowest perturbation achieved for the minimal amount of queries submitted. Prominent examples are HSJA [7], Guessing Smart (BAGS) [29], Sign-Opt [33], Policy-driven (PDA) [8], QEBA [34], and SurFree [35].
White-box attacks like C&W [17] cannot function in black-box environments as there is no closed-form description of the inference pipeline. To facilitate optimization, decision-based attacks commonly initialize from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation. This switch allows the task to be solved continuously, by minimizing the perturbation while always staying on the adversarial side of the boundary. Decision-based attacks share further common aspects in their functioning, ones that we can abstract through: given starting and original samples $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{g}$ and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{c}$ respectively, the goal is to iteratively propose adversarial candidates $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{t}$, until the distance $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\delta=d(x_{t},x_{c})$ is minimized. This process follows different algorithmic approaches that represent different geometrical intuitions; we can describe it more generally by means of a candidate generation policy:
@@
that given $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack, generates a candidate $x_{t}$. If we consider that attacks execute over discrete time steps, and assume the model always answers, then the attack procedure can be construed as a Markov Decision Process (MDP) to be solved, by finding the parameters $\theta$ that minimize $\delta$ in the least amount of queries.
Consider now a multinomial image classification model $\mathcal{M}$ under attack, with a discriminant function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes. By definition, black-box environments preclude access to these probabilities; instead one can only observe the decision of the classifier $C$ that returns the highest probability class:
@@
with $D$ being the decision function, here $D=\operatorname*{arg\,max}$. The goal in targeted attacks is to change the decision $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}c_{g}\in[m]$ for a correctly classified example $x$, to a predefined target class $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}c_{o}\neq c_{g}$. This process can be facilitated through a function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\psi$ which given a perturbed example $x_{t}$ at step $t$, it returns a binary indicator of success:
@@
As long as the model responds, $\psi$ can always be evaluated, it thus constitutes an essential mechanism upon which decision-based attacks build. The adversarial goal can then be described as the following constrained optimization problem:
@@
where the distance metric $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$. As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this optimization task highlights the fuzzy nature of adversarial examples, something that is further aggravated in domains where visual proximity is of little importance. Successful or unsuccessful adversarial examples are therefore delimited by an threshold $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$. While this constraint might not intuitively translate to non-visual domains, we nonetheless contend that minimal perturbation still remains a desideratum of adversarial examples.
The fact that real-world attacks are black-box and decision-based does not make them any less effective. For instance, HSJA is guaranteed to converge to a stationary point of Eq. (5). Given typical $\epsilon$ values for imperceptibility, this results in high attack success rates, even against adversarially trained models. The limited effectiveness of adversarial training against decision-based attacks can be attributed to the fundamentally out-of-distribution (OOD) nature of adversarial examples, and the inherent challenges in the saddle point optimization problem in Eq. (1) that make it difficult for algorithms to converge to a global solution. Additionally, decision-based attacks are challenging to incorporate during stochastic gradient descent (SGD) because these attacks rely on boundary navigation, and the farther the model is from convergence, the less effective the attack is. Scalability is another concern: while adversarial training typically involves a few steps (1-50) of white-box attacks like FGSM or PGD, decision-based attacks can take orders of magnitude more steps (queries) to produce an adversarial example.
Decision-based attacks search for the optimal parameters $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\theta^{*}$ of the generation policy (2), those that given $x^{i}_{c}$, with $i$ denoting the i-th adversarial episode, minimize Eq. (5) in expectation: